40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 46 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  343   Sun Sep 5 20:54:11 2010 taraNotesNoiseBudgetRIN from PMC and RCAV

 SInce the Faraday isolator alter the beam path and beam profile a bit. The mode matching efficiency drops.

I only realigned the beam to PMC and RCAV. Lenses before PMC haven't been moved yet.

                    efficiency (P_in/P_out)              DC level (from PD behind the cavity)

PMC        18.2mW/27.9mW = 65%                   2.5V

RCAV       4.56/4.94 [mW]  = 92% (this drops from ~96%)    1.3V

 

pmc_n and rc_n (blue/green) were taken together with Faraday isolator installed .

 

pmc and rc with no isolator (purple/brown) are data from Aug 30, 2010.    

 

pmc blocked (red)  is PMC RIN when the beam after the PMC is blocked, RCAV loop disabled, SLOWDC control is disabled.

There is no significant change here, I have to think about it.

  342   Sun Sep 5 19:41:45 2010 taraDailyProgress Faraday Isolator added

 Frank showed me that the RIN level from PMC is too high and caused by back reflection. 

PMC's psd is reduced by blocking the beam going to ACAV and RCAV.  This can be seen

on SR785, by blocking the beam.

So I removed PBS in front of the laser, and replaced it with a Faraday isolator. 

The isolator is temporarily mounted on a post, we will use a more rigid  V-shape block holder later.

The isolator shifts the beam path a bit, and the beam is needed to be re-align to the PMC. 

I couldn't finished it by Friday night, I think Frank re-aligned it already.

I'll measure RIN  and compare them again.

  341   Fri Sep 3 01:45:54 2010 taraSummaryLaserRIN comparison between PMC and RCAV

 When both cavities are locked, I measured

1) power fluctuation behind ACAV (DC@0.55V) and RCAV (DC@2.2V)

2) power fluctuation behind PMC (DC@3.17V)and RCAV (DC@2.2V)

the HEPA filter was on, the SLOWDC loop was engaged.

SR785 is set to AC couple mode, auto range on both channels.

 

The level of RIN from RCAV from two measurements are a bit different,see fig2.

RIN from ACAV is not significantly higher than that of RCAV. This is good news, I thought AOM might pose more pointing instability to the system.

  340   Thu Sep 2 20:31:01 2010 taraNotesComputersmedm screens are frozen

50 mins ago, medm on Sun machine is frozen, most channels are inactive (except  ACAV_HEATER, and it goes up to 7 which is the maximum value). 

I reboot the PSL crate, but ACAV_HEATER is still at 7 and heat up ACAV to 39 C.

So, I realized that I had to reboot another crate. Now ACAV_HEATER is at 4.6 and decreasing.

Now the system should try to bring the temperature back and might take many hours.

  339   Thu Sep 2 18:54:10 2010 taraNotesComputersACAV,RCAV's pds connected to DAQ

 

ACAV's PD for transmitted beam is connected to PSL1

RCAV's PD for transmitted beam is connected to PSL2

 

 

And now RCAV_RCTRANSPD dropped from ~5.3 to ~5.05

               ACAV_RCTRANSPD dropped from  ~2.5 to ~2.4

  338   Thu Sep 2 17:11:39 2010 taraNotesRefCavboth cavites locked

ACAV_TEMPAVG still adjusting it self to 37.2 (SETPT)  three overshoots in the last 8 hrs (+/-0.01 C)

 

                     ACAV                   RCAV

SETPT             37.2                  35

TEMPAVG    37.215               35.005

SLOWDC       N/A            -0.10932

VCOMON              1.11 V

 

*RCAV_TEMPAVG is very noisy compared to ACAV_TEMPAVG

  337   Wed Sep 1 15:22:39 2010 taraNotesRefCavcurrent parameters for cavities

   Temperature on both cavities are fairly constant over -8 hrs, still can't simultaneously lock both cavities.                        

                              ACAV                      RCAV

XXX_SETPT          37.3                       35.0

XXX_TEMPAVG        37.301                 34.997

SLOWDC              -0.11645 (VCOMON center @ 0),    -0.1114

XXX_HEATER      4.72971              (FSS)1.49154

 

Since everything is stable now, I'll lock RCAV and adjust ACAV temperature to lock them together withing AOM range.

I'll change ACAV_SETPT from 37.3->37.4, to see how SLOWDC has to change to lock ACAV.

 

Wed Sep 01 21:04:24 2010

After 6 hrs, ACAV_TEMPAVG is stable around 37.4 C.

SLOWDC for ACAV is -0.1200 (VCOMON center @ 0)

 This means For ACAV, dV(slow dc) per dT(ACAV) = (0.1203-0.11645)/0.1 = 0.039 V/K.

Since RCAV is locked when SLOWDC = -0.1114, we have to decrease the temperature from 37.4 C by (0.1203 - 0.1114)/0.039 = 0.2282 K.

That is 37.4 - 0.2282 = 37.1718.

Now I set ACAV_SETPT to 37.2 C.

  336   Wed Sep 1 00:27:34 2010 taraNotesBEATsymmetric optic layout for beat measurement

 A tentative layout for almost symmetric layout for beat measurement. The double 1/4 objects  in the beam paths should be 1/4 and 1/2 waveplates.

Lenses might be used to focus the beam somewhere.

  335   Tue Aug 31 21:59:19 2010 FrankSummaryComputersstill network problems

i tried to figure out where the network problems come from. Looks like it's the fiber connection between fb1 and the switch in the PSL lab.

Here a result from a simple ping between fb1 and the other computers. It acrually doesn't matter which one.

--- 10.0.0.1 ping statistics ---
1000 packets transmitted, 817 received, 18% packet loss, time 201397ms
rtt min/avg/max/mdev = 0.169/0.233/0.381/0.025 ms

--- 10.0.0.2 ping statistics ---
1000 packets transmitted, 786 received, 21% packet loss, time 202036ms
rtt min/avg/max/mdev = 0.617/0.694/2.658/0.152 ms

--- 10.0.0.3 ping statistics ---
1000 packets transmitted, 796 received, 20% packet loss, time 201696ms
rtt min/avg/max/mdev = 0.410/0.453/2.655/0.081 ms

Pings between computers within the PSL lab but connected to the same switch are OK:

--- 10.0.0.1 ping statistics ---
1000 packets transmitted, 1000 received, 0% packet loss, time 202998ms
rtt min/avg/max/mdev = 0.000/1.092/13.755/1.340 ms

--- 10.0.0.2 ping statistics ---
1000 packets transmitted, 1000 received, 0% packet loss, time 203161ms
rtt min/avg/max/mdev = 0.200/1.719/13.230/1.545 ms

So i think it's the fiber connection.

 

  334   Tue Aug 31 21:37:11 2010 FrankNotesRefCavchanged acav setpt back to 37.3

Quote:

don't change anything until it is stable and you thought about it carefully. The system was totally out of control, check the dataviewer!.
So there is no way to make an estimate for the values as you have no equilibrium!! That takes several hours after ANY change to the system, like rebooting the crate or turning on a temp control loop.
Remember: time constants are several hours!

So be patient and wait until tomorrow or late tonight before you make estimates and changes. Both temperatures are NOT close to their original setpoint and largely fluctuating.
RCAV is at 34.8 instead of 35 and ACAV changed from 37.1 to 37.45 within the last hours.
So no way to make the right adjustments right now, as your laser SLOWDC value does not reflect the actual temperature of the chamber.

Changed it back to the original settings (37.3) as those were working over several days before the system stopped working properly...

Quote:

I adjust SLOWDC to see what value to lock each cavity.

For ACAV I adjust SLOWDC so that VCOMON fluctuates around 0 and record the value of SLOWDC

I adjust ACAV_SETPT to see how the DC changes (in RED), it seems I have to reduce the ACAV_SETPT

                             ACAV                              RCAV

SLOWDC       -0.1118 ->-0.1151             -0.1051

SETPT           37.3      -> 37.35                       35

 

Now, ACAV_SETPT is changed to 37.15, but ACAV_TEMPAVG goes up (after 30mins.) to ~37.43 (from 37.3).

The variance from ACAV's 4 temp sensors are getting larger.

 

 Frank, did you remotely change ACAV_SETPT back to 37.3? I came back and it was back to 37.3.

  333   Tue Aug 31 18:33:13 2010 FrankNotesRefCavchanged acav setpt back to 37.3

don't change anything until it is stable and you thought about it carefully. The system was totally out of control, check the dataviewer!.
So there is no way to make an estimate for the values as you have no equilibrium!! That takes several hours after ANY change to the system, like rebooting the crate or turning on a temp control loop.
Remember: time constants are several hours!

So be patient and wait until tomorrow or late tonight before you make estimates and changes. Both temperatures are NOT close to their original setpoint and largely fluctuating.
RCAV is at 34.8 instead of 35 and ACAV changed from 37.1 to 37.45 within the last hours.
So no way to make the right adjustments right now, as your laser SLOWDC value does not reflect the actual temperature of the chamber.

Changed it back to the original settings (37.3) as those were working over several days before the system stopped working properly...

Quote:

I adjust SLOWDC to see what value to lock each cavity.

For ACAV I adjust SLOWDC so that VCOMON fluctuates around 0 and record the value of SLOWDC

I adjust ACAV_SETPT to see how the DC changes (in RED), it seems I have to reduce the ACAV_SETPT

                             ACAV                              RCAV

SLOWDC       -0.1118 ->-0.1151             -0.1051

SETPT           37.3      -> 37.35                       35

 

Now, ACAV_SETPT is changed to 37.15, but ACAV_TEMPAVG goes up (after 30mins.) to ~37.43 (from 37.3).

The variance from ACAV's 4 temp sensors are getting larger.

 

  332   Tue Aug 31 18:20:42 2010 taraNotesRefCavchange acav setpt

I adjust SLOWDC to see what value to lock each cavity.

For ACAV I adjust SLOWDC so that VCOMON fluctuates around 0 and record the value of SLOWDC

I adjust ACAV_SETPT to see how the DC changes (in RED), it seems I have to reduce the ACAV_SETPT

                             ACAV                              RCAV

SLOWDC       -0.1118 ->-0.1151             -0.1051

SETPT           37.3      -> 37.35                       35

 

Now, ACAV_SETPT is changed to 37.15, but ACAV_TEMPAVG goes up (after 30mins.) to ~37.43 (from 37.3).

The variance from ACAV's 4 temp sensors are getting larger.

  331   Tue Aug 31 16:42:04 2010 taraNotesDAQC3:PSL-ACAV_RCTRANSPD broken?

Quote:

Quote:

was it grey? If yes the cable already connected to the T is the cable going to the DAQ. That's why there is a T already. Don't disconnect any signals without checking in advance what they are and what they are used for. Some of them are required for locking, e.g. the transmitted light is required for the servo as a feedback if it is locked or not for the temp feedback. If you have to disconnect anything do it but put an entry in the elog.

Quote:

Quote:

can you plz check this channel as it is constantly at -14.4797 since yesterday. Did you disconnect it? If yes, plz elog those things!! Everything must be eloged

 I didn't do anything on that channel. I just removed the BNC cable that connect to a T behind the PD and connected it to an oscilloscope this afternoon. I'll take a look.

 

 No, it's the black one that goes to a multimeter. The grey cable to DAQ is untouched.

 I checked the beam. It hit the PD and gives ~500 mV output when only ACAV is locked. when I disconnected the grey cable that connects the PD to DAQ, the

value is still the same (-14.5).

The connection at DAQ looks good.

The cable is good, the signal can go through.

I thought 0the setup in  RCAV.db and ACAV.db files were wrong so I changed it to the red values

                             rcav.db                     acav.db

field DTYP        VMIVME-3113          VMIVME-3123 ->VMIVME-3113

INP                   #c0 S61@                     #c0 S1@  -> #c0 S62@

 ,and now it's working.

I remember changing these before Frank went to Hanford, I might not save the file. Anyway, now I saved it.

The fluctuation in VCOMON goes down to less than +/- 1V.

  330   Tue Aug 31 16:15:51 2010 taraNotesComputersdead channel, C3:PSL-ACAV_TEMPAVG

Quote:

be carefull with the data you are taking right now. it's wrong for your projection as the power fluctuations are different when locking only the ACAV using the AOM. The largest contributor might be the pointing from the AOM itself, which is different if the laser isn't locked to the other cavityat the same time.

Why don't you use the new fast channels you have hooked up last week? And don't forget to change the names of those :-)

Quote:

Quote:

looks like some fault of the database. /usr1/epics/psl/db/acav.db does not contain the correct entry. check rcav.db and copy the record for "C3:PSL-RCAV_TEMPAVG" into acav.db. Then simply change "RCAV" into "ACAV" everywhere for this record. Also change the setpoint for the ACAV in the startup.cmd file to 37.3 and the ACAV-heater value to 4.653. Those are the latest values when both where locked for several hours. Reset the crate again.

Quote:

Yesterday, I reset the PSL crate behind the SUN computer, but the channel C3:PSL-ACAV_TEMPAVG is stil inactive.

 

 

 Will do in a moment, I'm taking data from ACAV for now just to compare with yesterday results.

 I didn't disabled the loop when I measured it, and yeah it looks bad. I think I'll just try to lock both cavities for now.

I'll connect the signals from PDs behind Rcav and Acav to the new fast channels(32k).

About changing the channels' names, I asked DMASS to help and he suggested to use them as they are for now because of the risk of screwing the system up from a typo.

  329   Tue Aug 31 16:04:30 2010 taraNotesDAQC3:PSL-ACAV_RCTRANSPD broken?

Quote:

was it grey? If yes the cable already connected to the T is the cable going to the DAQ. That's why there is a T already. Don't disconnect any signals without checking in advance what they are and what they are used for. Some of them are required for locking, e.g. the transmitted light is required for the servo as a feedback if it is locked or not for the temp feedback. If you have to disconnect anything do it but put an entry in the elog.

Quote:

Quote:

can you plz check this channel as it is constantly at -14.4797 since yesterday. Did you disconnect it? If yes, plz elog those things!! Everything must be eloged

 I didn't do anything on that channel. I just removed the BNC cable that connect to a T behind the PD and connected it to an oscilloscope this afternoon. I'll take a look.

 

 No, it's the black one that goes to a multimeter. The grey cable to DAQ is untouched.

  328   Tue Aug 31 15:52:43 2010 FrankNotesDAQC3:PSL-ACAV_RCTRANSPD broken?

was it grey? If yes the cable already connected to the T is the cable going to the DAQ. That's why there is a T already. Don't disconnect any signals without checking in advance what they are and what they are used for. Some of them are required for locking, e.g. the transmitted light is required for the servo as a feedback if it is locked or not for the temp feedback. If you have to disconnect anything do it but put an entry in the elog.

Quote:

Quote:

can you plz check this channel as it is constantly at -14.4797 since yesterday. Did you disconnect it? If yes, plz elog those things!! Everything must be eloged

 I didn't do anything on that channel. I just removed the BNC cable that connect to a T behind the PD and connected it to an oscilloscope this afternoon. I'll take a look.

 

  327   Tue Aug 31 15:46:13 2010 taraNotesDAQC3:PSL-ACAV_RCTRANSPD broken?

Quote:

can you plz check this channel as it is constantly at -14.4797 since yesterday. Did you disconnect it? If yes, plz elog those things!! Everything must be eloged

 I didn't do anything on that channel. I just removed the BNC cable that connect to a T behind the PD and connected it to an oscilloscope this afternoon. I'll take a look.

  326   Tue Aug 31 15:15:44 2010 FrankNotesDAQhow to measure ACAV channels without RCAV locked

if you wanna measure anything in the ACAV loop without the RCAV locked, don't forget to DISABLE the FSS loop.
If you don't, your measurements will be shit as the PC is fluctuating like hell making the laser noisy which can't be handled by the other loops due to limited bandwidth...

  325   Tue Aug 31 15:11:28 2010 FrankNotesDAQC3:PSL-ACAV_RCTRANSPD broken?

can you plz check this channel as it is constantly at -14.4797 since yesterday. Did you disconnect it? If yes, plz elog those things!! Everything must be eloged

  324   Tue Aug 31 15:04:51 2010 FrankNotesComputersdead channel, C3:PSL-ACAV_TEMPAVG

be carefull with the data you are taking right now. it's wrong for your projection as the power fluctuations are different when locking only the ACAV using the AOM. The largest contributor might be the pointing from the AOM itself, which is different if the laser isn't locked to the other cavityat the same time.

Why don't you use the new fast channels you have hooked up last week? And don't forget to change the names of those :-)

Quote:

Quote:

looks like some fault of the database. /usr1/epics/psl/db/acav.db does not contain the correct entry. check rcav.db and copy the record for "C3:PSL-RCAV_TEMPAVG" into acav.db. Then simply change "RCAV" into "ACAV" everywhere for this record. Also change the setpoint for the ACAV in the startup.cmd file to 37.3 and the ACAV-heater value to 4.653. Those are the latest values when both where locked for several hours. Reset the crate again.

Quote:

Yesterday, I reset the PSL crate behind the SUN computer, but the channel C3:PSL-ACAV_TEMPAVG is stil inactive.

 

 

 Will do in a moment, I'm taking data from ACAV for now just to compare with yesterday results.

  323   Tue Aug 31 14:52:41 2010 FrankNotesComputersdead channel, C3:PSL-ACAV_TEMPAVG

made the changes a minute ago. simply reboot after changing the values in the startup.cmd (those i didn't change)

Quote:

Quote:

looks like some fault of the database. /usr1/epics/psl/db/acav.db does not contain the correct entry. check rcav.db and copy the record for "C3:PSL-RCAV_TEMPAVG" into acav.db. Then simply change "RCAV" into "ACAV" everywhere for this record. Also change the setpoint for the ACAV in the startup.cmd file to 37.3 and the ACAV-heater value to 4.653. Those are the latest values when both where locked for several hours. Reset the crate again.

Quote:

Yesterday, I reset the PSL crate behind the SUN computer, but the channel C3:PSL-ACAV_TEMPAVG is stil inactive.

 

 

 Will do in a moment, I'm taking data from ACAV for now just to compare with yesterday results.

 

  322   Tue Aug 31 14:50:53 2010 taraNotesComputersdead channel, C3:PSL-ACAV_TEMPAVG

Quote:

looks like some fault of the database. /usr1/epics/psl/db/acav.db does not contain the correct entry. check rcav.db and copy the record for "C3:PSL-RCAV_TEMPAVG" into acav.db. Then simply change "RCAV" into "ACAV" everywhere for this record. Also change the setpoint for the ACAV in the startup.cmd file to 37.3 and the ACAV-heater value to 4.653. Those are the latest values when both where locked for several hours. Reset the crate again.

Quote:

Yesterday, I reset the PSL crate behind the SUN computer, but the channel C3:PSL-ACAV_TEMPAVG is stil inactive.

 

 

 Will do in a moment, I'm taking data from ACAV for now just to compare with yesterday results.

  321   Tue Aug 31 13:40:00 2010 FrankNotesComputersdead channel, C3:PSL-ACAV_TEMPAVG

looks like some fault of the database. /usr1/epics/psl/db/acav.db does not contain the correct entry. check rcav.db and copy the record for "C3:PSL-RCAV_TEMPAVG" into acav.db. Then simply change "RCAV" into "ACAV" everywhere for this record. Also change the setpoint for the ACAV in the startup.cmd file to 37.3 and the ACAV-heater value to 4.653. Those are the latest values when both where locked for several hours. Reset the crate again.

Quote:

Yesterday, I reset the PSL crate behind the SUN computer, but the channel C3:PSL-ACAV_TEMPAVG is stil inactive.

 

 

  320   Tue Aug 31 13:33:02 2010 taraNotesComputersdead channel, C3:PSL-ACAV_TEMPAVG

Yesterday, I reset the PSL crate behind the SUN computer, but the channel C3:PSL-ACAV_TEMPAVG is stil inactive.

 

  319   Tue Aug 31 13:26:32 2010 FrankNotesEnvironmentantack (ants attack)

any results from the poison from 40m they tried in the TCS lab? Did it work?If yes we should get some for the other labs too....

Quote:

 I think ants are now building their colony in the lab, even though there is no garbage, they can find some food around here.

That seems like a cricket's leg (see attached pic). A few of them walk around the optic table/ optics too.

I think we should let an aardvark roam around the lab for a day.

 

 

  318   Mon Aug 30 23:47:30 2010 taraNotesEnvironmentantack (ants attack)

 I think ants are now building their colony in the lab, even though there is no garbage, they can find some food around here.

That seems like a cricket's leg (see attached pic). A few of them walk around the optic table/ optics too.

I think we should let an aardvark roam around the lab for a day.

 

  317   Mon Aug 30 23:25:06 2010 taraSummaryRefCavRIN comparison between PMC and RCAV

 I measured RIN from beam behind PMC and RCAV. Photodiodes are connected to SR785 and measured at the same time. 

DC level from PMC and RCAV are 3.56 V and 1.58 V, respectively.

The HEPA filters above the table were turned off during one measurement to see how much it would effect the RIN (turn out to be not that much)

Only PMC and RCAV were locked during the measurement.

The results are plotted together with RIN from 40m's RCAV ( 40m elog. ) PSL has lower noise at lower frequency f< 50Hz, and about 1.5 order of magnitude higher

between 50Hz to 10 kHz. 

I haven't tried to optimize any cavities yet. This is just walk-in,plug-in and measure as a reference for the current setup.

if I try to re-center the beam into the cavities, the result might be better (or worse).

  316   Mon Aug 30 22:16:51 2010 FrankNotesComputersnote for resetting psl crate

set temp for the ACAV should be 37.3, that was the value since about a week. see here : http://131.215.115.52:8080/PSL_Lab/307

Quote:

  C3:PSL-ACAV_TEMPAVG that tells the average temperature from ACAV is dead. Frank told me to reboot the PSL crate.

                                       RCAV      ACAV

C3:PSL-XXXX_SETPT      35.0       36.75

C3:PSL-XXX_HEATER    1.47521   4.9

 

I rebooted the crate, reload the cavity.stp file, the channel is still inactive. (only this channel)

 

  315   Mon Aug 30 20:11:51 2010 taraNotesComputersnote for resetting psl crate

  C3:PSL-ACAV_TEMPAVG that tells the average temperature from ACAV is dead. Frank told me to reboot the PSL crate.

                                       RCAV      ACAV

C3:PSL-XXXX_SETPT      35.0       36.75

C3:PSL-XXX_HEATER    1.47521   4.9

 

I rebooted the crate, reload the cavity.stp file, the channel is still inactive. (only this channel)

  314   Mon Aug 30 11:30:22 2010 taraNotesComputersmedm screens are frozen

I was measuring the power fluctuation behind PMC and RCAV, during that the beam was not locked and the medm screens were frozen, so I restarted the crate.

  313   Fri Aug 27 20:08:56 2010 taracNotesNoiseBudgetCOMSOL

  I use COMSOL to simulate the temperature distribution inside the mirror. 

The mirrors are separate into 3 regions for different meshing quality see attached figures.

This model assume 0.95 emissivity on the surface, and 310 K at the surface between the mirror and the spacer.

The temperature between 10 mW and 10mW + RIN*10mW is plotted and fitted with polynomial function of order 9.

*the unit of y axis on fig3 is Kelvin

The Temperature difference will be applied to the coating surface and calculated the upper limit for photo thermal noise. 

The function is

 

T = p1*z.^9 + p2*z.^8 + p3*z.^7 + p4*z.^6 + 

    p5*z.^5 + p6*z.^4 + p7*z.^3 +

    p8*z.^2 + p9*z + p10;

 
where

 

 

       p1 =  -5.335e-10 ; %(-5.806e-10, -4.864e-10)

       p2 =   1.602e-08 ;% (1.474e-08, 1.731e-08)

       p3 =  -2.041e-07 ;% (-2.188e-07, -1.894e-07)

       p4 =   1.438e-06 ;% (1.347e-06, 1.528e-06)

       p5 =  -6.144e-06  ;%(-6.47e-06, -5.818e-06)

       p6 =   1.645e-05  ;%(1.576e-05, 1.715e-05)

       p7 =  -2.773e-05  ;%(-2.858e-05, -2.687e-05)

       p8 =   2.912e-05  ;%(2.857e-05, 2.967e-05)

       p9 =  -1.893e-05  ;%(-1.908e-05, -1.877e-05)

       p10 =   7.548e-06 ;% (7.534e-06, 7.562e-06)

 

 

 

  312   Fri Aug 27 19:02:24 2010 FrankNotesComputersDAQ

plz change the channel names. Those names are bullshit...

Quote:

Dmass helps me initializing two channels for DAQ.

PSL1 connected to channel 28 which is C2:ATF-ACCoup_AC2_OUT_DAQ

PSL2 to channel 29 which is                    C2:ATF-ACCoup_AC3_OUT_DAQ

 

  311   Fri Aug 27 18:49:43 2010 taraNotesBEATboth cavities relocked

 

rebooted the crate.

Quote:

relocked both cavities. The second crate containing the output card to control the ACAV power supply stopped working and i can't login from here. It's blocking the telnet request.

So Tara, plz reboot the crate in the other "rack", above the SR785. Simply push the reset button... Thanks

 

  310   Fri Aug 27 18:39:23 2010 taraNotesComputersDAQ

Dmass helps me initializing two channels for DAQ.

PSL1 connected to channel 28 which is C2:ATF-ACCoup_AC2_OUT_DAQ

PSL2 to channel 29 which is                    C2:ATF-ACCoup_AC3_OUT_DAQ

  309   Thu Aug 26 23:18:12 2010 FrankNotesBEATboth cavities relocked

relocked both cavities. The second crate containing the output card to control the ACAV power supply stopped working and i can't login from here. It's blocking the telnet request.

So Tara, plz reboot the crate in the other "rack", above the SR785. Simply push the reset button... Thanks

  308   Thu Aug 26 23:00:35 2010 FrankNotesBEATlock-loss of both cavities

i lost lock of both cavities for some unknown reason. Also the crate seemed to stop working as some channels are dead now...

  307   Thu Aug 26 19:58:06 2010 FrankNotesBEATchanged ACAV setpoint temperature - both cavities locked

changed the ACAV temp to 37.3 degrees. Both cavities are locked now.

I'll monitor the VCO input signal to center the range by adjusting the temp a little bit more.

  306   Mon Aug 23 21:50:47 2010 FrankNotesRefCavchanged ACAV setpoint temperature

changed setpoint for ACAV to 37.2 degrees, disabled the FSS loop and enabled feedback to laser temp using VCO input signal (C3:PSL-ACAV_VCOMON)

RCAV resonant at -0.1154,  AOM centered at -0.0987

  305   Fri Aug 20 21:03:16 2010 FrankSummaryDAQchannel list for fb1 updated and framebuilder restarted

i updated the database for epics channels for the PSL lab and restarted the framebuilder on fb1.
I didn't restart the framebuilder on fb0, but once it will be restarted it' will load the new channel list too.

I saw that something seems to be wrong with the network as we have several lost packets which also shows up in the saved data.
I will investigate this later as it seems to be a problem for the slow channels from the PSL lab only.

  304   Fri Aug 20 18:59:40 2010 MeganPhotosEnvironmentTemperature Stabilization?

To test the temperature stability of our cavities, Frank and I decided to run the heat gun at 800 degrees Fahrenheit for a bit of time. Coincidentally, this is the perfect temperature for roasting marshmallows. Complete coincidence, of course. After setting up a stable marshmallow roaster, er, heat gun,

IMG_4914.jpg

we roasted marshmallows successfully (with the help of Zach) for a little while.

IMG_4913.jpgIMG_4915.jpg

While this was a successful test for our tastebuds and need for sugar, the impact on the cavities is uncertain, and so this entire setup should probably be tested again. Including the marshmallows.

  303   Fri Aug 20 00:54:48 2010 FrankSummaryDAQPSL RT fronend code shut down

i've killed the PSL RT frontend. Everything else should be ok, so plz check if everything is working.

  302   Fri Aug 20 00:14:09 2010 FrankSummaryDAQchanges at VCO connecting block at x-connect

i replaced some cables going from the 9-pin D-SUB connection block connecting the VCO with the DAQ by bi-colored, twisted ones.
Someone used yellow cables for everything when he set this up. Now the ADC and DAC channels are bi-colored (black-LO, yellow-HI).

I also removed one connection from ground to the rack frame, so protective earth. I don't see why we need that as we have one at the power supply already. There ground=earth.

  301   Fri Aug 20 00:04:14 2010 FrankSummaryDAQchannels from PSL RT system

i disconnected all signal from the PSL RT system running on fb0 this afternoon, so this model doesn't have to run anymore.
They are all hooked up to the VME system now.

I will remove the configuration file for those channels from the fb0 framebuilder the next day. Right now it's still running but not used anymore...

  300   Fri Aug 20 00:00:42 2010 FrankSummaryDAQnew/old channels for transmitted light of refcav

as we moved the channels from the 16bit card to the 12bit card we also changed the channel name for the transmitted light of the reference cavity.
The original (historic) channel name was C3:PSL-FSS_RCTRANSPD, but since we have two cavities now we started to separate both systems by using the subsystem names RCAV and ACAV quite some time ago.

So far we still use some of the old channel names like the ones for the VCO as "FSS" channels, even if they belong to the other cavity since some while.
As it is quite a lot of work to change everything (you have to change all screens, scripts etc too) we didn't change those yet.

As we moved the PD channels today, we renamed the C3:PSL-FSS_RCTRANSPD into C3:PSL-RCAV_RCTRANSPD.
So plz use the new name from now on. We created a calc record entry for the old channel name which basically copies the value from the new, physical channel.
So all the old screens and scripts are still working. Once we changed all things and if we find the time we disable that channel and check if everything is still working.
Right now i don't see a problem in having both in parallel and use the new channel name wherever me make changes....

  299   Thu Aug 19 19:19:26 2010 taraNotesComputersslow actuator note

ACAV 0.2602 V

RCAV 0.1102 V

  298   Thu Aug 19 19:17:44 2010 FrankSummaryDAQremoved connecting blocks J21 & J22

removed both 9-pin D-SUB blocks (blocks J21 & J22) which were used for the RF photodiodes (PMC, RCAV) before (only to power them).
As we hooked them up to individual power supplies a long time ago they are free to use for the temp readout of the cavities.

  297   Thu Aug 19 18:56:29 2010 Megan, FrankSummaryDAQmoved PD channel to 12bit card (VMIC 3113A)

moved PD channel to 12bit card (VMIC 3113A)

C3:PSL-PMC_PMCTRANSPD now connected to CH63, pins 64(HI)/63(LO) on block J3 (J3-3113A-P3)

C3:PSL-ACAV_RCTRANSPD now connected to CH62, pins 62(HI)/61(LO) on block J3 (J3-3113A-P3)

C3:PSL-RCAV_RCTRANSPD now connected to CH61, pins 60(HI)/59(LO) on block J3 (J3-3113A-P3)

  296   Wed Aug 18 20:27:00 2010 FrankNotesDAQVMIC3123 address jumper settings for 1st card in PSL crate

jumper.jpg

  295   Wed Aug 18 20:16:24 2010 FrankNotesDAQtemp sensor 2 of RCAV is too noisy

temp sensor number 2 of the reference cavity shows much more (horrible) noise compared to the other ones (2K pkpk).

broken, oscillating?

Will check this tomorrow.

 

- personal notes -

things to do: change limit of both heaters to match new situation. much less power required and messing up things causes both cavities to heat up too much (50 C)...

  294   Wed Aug 18 20:11:44 2010 FrankSummaryDAQsome signals moved to VME DAQ system

moved the first 8 signals, the 4 individual sensors for each cavity, to the VME DAQ system.
Because i would have to open the foam box in order change the cables, i extended the existing ones for now to see if everything is working.
Once all channels are configured and everything is working i will open the box and make all the changes.
I also have to add the voltage reference for the second board and add the averaged channel from that board to the DAQ as well.
I need two more of the 9-pin D-SUB connector blocks or i have to make breakout cables instead.

Also added the VCO input monitor signal to the DAQ: channel name is C3:PSL-ACAV_VCOMON.
This channel can be used to keep the VCO signal centered in the range by feeding back to the temp setpoint of the cavity or the laser temp if locked alone (to match resonance of both cavities)

ELOG V3.1.3-